# Topological normal generation of big mapping class groups

### v2

@@ 1. Introduction
Let $S$ be an orientable surface. We say that $S$ is of finite type if its fundamental group is finitely generated. Otherwise, we will call $S$ is called a surface of infinite type. Equivalently, $S$ has infinite type if it has either infinitely many punctures, infinitely many genus, or both. A mapping class group of an infinite type surface $S$ is usually called a big mapping class group. Throughout this article we denote the big mapping class group of $S$ by $\mathrm{Map}(S)$. For a more detailed study of the big mapping class groups, we refer to the reader to [AV20].
@@ 1.2. First cohomology classes $H^{1}(\mathrm{FMap}(S),\mathbb{Z})$
The maximality of an end is derived from the partial order on the end space introduced in [MR23]. Mann and Rafi proved that under this partial order, the maximal element exists. Choose an immediate predecessor $z$ which both accumulates to maximal ends $A$ and $B$. (For precise definition, we refer Section 2.) We number each $z$ by integers in a sequence so that $\{z_{n}\}_{n\in\mathbb{Z}}$ accumulates to $A$ (or $B$) when $n$ goes to $\infty$ (or $\infty$- ∞, respectively). The generalized shift $\eta_{A,B,z}$ is a shift map which sends $z$ of index $n$ to $z$ with index $n+1$. In simple terms, the cohomology class we investigate counts how many $z$’s are shifted. Collecting all such countings (including handle shifts), we have the following structure for $\mathrm{FMap}(S)$.
@@ 1.3. Normal generating set of big mapping class groups
There have been numerous studies on the (normal) generation of the big mapping class group, as finding a nice generating set for a given group is crucial for understanding its structure and properties. In [APV20], the authors prove that the set of Dehn twists topologically normally generates $\mathrm{Map}(S)$ if and only if $S$ has at most one non-planar end. In [CC22], the authors show that for the finite type surface $S$ with a Cantor set $K$ embedded, $\mathrm{Map}(S-K)$ is normally generated by torsion normal generator of $\mathrm{Map}(S)$ unless $S$ has genus $2$ with $45k+$ 4 punctures for some $k\geq 0$. In [HHHM+22], the authors studied topological properties of conjucagy class in $\mathrm{Map}(S)$. They showed that all conjugacy classes in $\mathrm{Map}(S)$ are meager and provided a condition for $\mathrm{Map}(S)$ to contain a dense conjugacy class. Additionally, as mentioned earlier, Vlamis ([Vla23], [Vla24]) proved that if $S$ is perfectly self-similar or telescoping, then $\mathrm{Map}(S)$ has a normal generator. In [ACCL21], the authors show that If $S$ is an infinite-genus surface, then $\mathrm{PMap}(S)$ is topologically generated by handle shifts. Also, Mann and Rafi [MR23] give a necessary and sufficient condition for a tame surface $S$ to admit a coarsely bounded generating set of $\mathrm{Map}(S)$ and explicitly constructed the CB generating set.
@@
Also, let $C:=\max_{A,B\in\mathcal{M}(S)}|E_{cp}(A,B)|$. Then $\mathrm{Map}(S)$ is topologically normally generated by $M(M+C-1)$ number of generators.
@@ 1.4. Bounds of the rank of the abelianization of $\mathrm{Map}(S)$
    The commutator subgroup $[\mathrm{Map}(S),\mathrm{Map}(S)]$ is Polish and both open and closed subgroup of $\mathrm{Map}(S)$.
@@
Corollary 8.10 Suppose $S$ is a finite connected sum of perfectly self-similar surfaces. Then the abelianization of $\mathrm{Map}(S)$ is finitely generated by at most $M(M+C-1)$.
@@ 2.2. Mann-Rafi’s classification
    $E(z)\cap(A-\{x_{A}\})\neq\emptyset\text{\quad and \quad}E(z)\cap(B-\{x_{B}\})\neq\emptyset$
@@ 3. Obstruction
[…]
@@
$f\mapsto(\text{ \# punctures in }\Sigma_{\gamma}-\text{\# punctures in }\Sigma_{f(\gamma)},\text{\# genus in }\Sigma_{\gamma}-\text{\# genus in }\Sigma_{f(\gamma)})$
@@ 4. Various shifts
×$\mathcal{B}\cong\mathbb{R}\times[-1,1]-\left(\bigcup_{m\in\mathbb{Z}}D_{\frac{1}{2}}(m,0)\right)$
@@
We label $\partial_{m}$ to each boundary of the deleted disk centered at $(m,0)$. We also label $\partial_{\pm}$ as long boundaries, ×{±$\mathbb{R}\times\{\pm 1\}$.
@@
We index each $z$ so that $z_{k}$ accumulates to one maximal end $A$ as $k$ goes to $\infty$- ∞, The other maximal end $B$ is accumulated by $z_{k}$’s with positive indices.
@@ 5. Cohomology classes of $\mathrm{FMap}(S)$
By construction, $\Phi_{z}(\eta_{A,B,z^{\prime}})=1$ if and only if $z\sim z^{\prime}$. In other words, $\Phi_{z}$ detects how many shifts of $z$ occur between $A,B$. This detection also works for other types of shift maps.(recall Definition 4.4 that there are $3$ types of generalized shifts.) Proposition 4.5 implies that even if the shift is not full, it differs only by the surface homeomorphism $T$ to the full shift. Since $T$ is generated by half twists (or $1/n$ Dehn twists), $\Phi_{z}(T)$ is equal to 00 by (3).
@@
Consider the standard basis $\{e_{i}\}$ of $\mathbb{Z}^{N-1}$, where $z\in E_{cp}(A_{0},A_{i})$ is an end corresponding to the $i$th entry. Let $\eta_{i}:=\eta_{A_{0},A_{i},z}$ Then $\Theta(\eta_{i})=e_{i}$. Since all $\eta_{i}$’s commute each other, the sequence splits. ∎
@@ 6. From $\mathrm{FMap}(S)$ to $\mathrm{Map}(S)$
    $\mathcal{P}$ measures the parity of the permutation on $\{x_{1},\cdots,x_{n}\}$, which sends odd permutations to $1$ and even permutations to 00, induced by the forgetful map $\mathrm{Map}(S)\twoheadrightarrow S_{n}\twoheadrightarrow\mathbb{Z}/2\mathbb{Z}$.
@@
    Let $i,j$ be distinct integers in $\{1,\cdots,n\}$. Choose $z\in Im(x)$ and let $\eta$ be a shift map of $z$ from $x_{i}$ to $x_{j}$. If $Im(x)$ only contains a handle, then $x$ must be a type of ends which is accumulated only by genus and take the shift map $\eta$ as a handle shift from $x_{i}$ to $x_{j}$. Recall that $\Theta_{z}:\mathrm{FMap}(S)\to\mathbb{Z}^{n-1}$. Define $\widetilde{\Theta_{z}}:\mathrm{FMap}(S)\to\mathbb{Z}^{n-1}\to\mathbb{Z}\to\mathbb{Z}/2\mathbb{Z}$, where the second map is the sum of all entries in $(n-1)$ tuple and the last map is a usual quotient map by $2\mathbb{Z}$.
@@
Let $f\in\mathrm{Map}(S)$. We first show that $\widetilde{\Theta_{z}}$ is invariant under conjugation. Conjugation by an elements in $\mathrm{FMap}(S)$ does not change the value, by its definition. Hence, the only conjugation which may invoke any differences is by half twists. Conjugation by half twists differs value to either permute entries in the $(n-1)$ tuple, or change signs of some entries. Therefore, taking the sum and quotient by $2\mathbb{Z}$ is invariant.
@@ 7.1. Half-spaces
We first decompose the surface $S$ as follows. Choose a subsurface $K$ embedded in $S$ which is homeomorphic to a sphere with $4$ boundaries, such that $S-K$ has $4$ connected components consisting of $S_{u}$ deleted with an open disk and $3$ homeomorphic copies of $S_{p}$ deleted with an open disk. By abusing the notation, we label each subsurface with $S_{u}$ and $S_{1},S_{2},S_{3}$, respectively. Note that the boundary of $S_{u}$ separates the $S_{u}$ part from the $S_{p}$ part. See Figure 7 also.
Figure 7. The surface decomposition by the $2$-sphere with $4$ boundaries at the center. There is a $S_{u}$ on the top, and the other $3$ subsurfaces are $S_{1},S_{2},S_{3}$, all homeomorphic to each other.
@@
Before we continue the argument, we first introduce which element has a dense conjugacy class. We follow the notation in [LV23a]. Let $S$ be uniquely self-similar surface and $x$ be a maximal end of $S$.
Proposition 7.8 (Proposition 4.2 in [LV23a]).
Let $c$ be a separating curve of $S$ and $\Omega_{c}$ denote the component of $S-c$ such that $x\in E(\Omega_{s})$. Suppose $G$ be a subgroup of $\mathrm{Map}(S)$ as follows.
@@ 8. Normal generating sets
Let $G$ be a subgroup of $\mathrm{Map}(K,\partial K)$ whose elements induce the mapping class in $\mathrm{Map}(S)$, except the Dehn twists along the boundary of $K$. Then $G$ is normally generated by $M(M-2)$ number of elements.
@@
Suppose $\gamma$ is a separating curve that intersects $2$ boundaries of $L$ and one of the connected components of $K-\gamma$ is a pair of pants. See also Figure 8. The conjugacy class of $T_{\gamma}$ depends only on which boundaries $\gamma$ meets. This is because, if we have another $\gamma^{\prime}$, then there are appropriate half twists $g,g^{\prime}$ supported in $K-\text{int}(L)$ so that $gg^{\prime}(\gamma)=\gamma^{\prime}$. Therefore we only need $\frac{M(M-1)}{2}$ curves for each pair of connected components of $K-\text{int}(L)$ Also if we choose another curve intersecting $2$ boundaries of $L$, then the Dehn twist along the curve can be generated by the lantern relation. Again by the lantern relation, the Dehn twist along any boundary of $L$ is generated by the half twists which are supported on the component of $K-\text{int}(L)$ which contains the boundary.
@@
Also, let $C:=\max_{A,B\in\mathcal{M}(S)}|E_{cp}(A,B)|$. Then $\mathrm{Map}(S)=\overline{\langle\langle\mathcal{S}\cup\mathcal{D}\cup\mathcal{H}\rangle\rangle}$. i.e., $\mathrm{Map}(S)$ is topologically normally generated by $M(M+C-1)$.
@@
Bounds on the minimal number of topological normal generators gives a bound of rank of its abelianization. By Field-Patel-Rasmussen [FPR22], they proved that the commutator subgroup of the mapping class group of surface $S-F$ we questioned above is closed is finitely generated and discrete with the quotient topology.
@@
    The commutator subgroup $[\mathrm{Map}(S),\mathrm{Map}(S)]$ is Polish and both open and closed subgroup of $\mathrm{Map}(S)$.
@@
Suppose $S$ is a finite connected sum of perfectly self-similar surfaces, possibly with finitely many punctures. Then the abelianization of $\mathrm{Map}(S)$ is finitely generated by at most $M(M+C-1)$.
@@ 9. Remarks on locally CB
In [LV23b], the authors proved that for an infinite type surface $S$, $\mathrm{Map}(S)$ has the virtual Rokhlin property (i.e., $\mathrm{Map}(S)$ contains a finite index subgroup which has the Rokhlin property) if and only if $S$ is spacious. More precisely, they found an open subgroup $\Gamma(S)$ which is an open, finite-index subgroup of $S$ with the Rokhlin property. We first recall the some definitions in [LV23b].
@@
    Suppose $S$ is a weakly self-similar and it is homeomorphic to $N$ or $N\#N$ for some uniquely self-similar surface $N$. Then $\mu\in E(S)$ is maximally stable if there is a nested subsequence $\Sigma_{n},n\in\mathbb{Z}_{\geq 0}$ such that $\Sigma_{n}$ is homeomorphic to $N$ with the open disk removed, $\Sigma_{n}$ is a neighborhood of $\mu$, and $\cap_{n\in\mathbb{Z}_{\geq 0}}$ is empty. Let $\Gamma(S)$ be a subgroup of $\operatorname{Homeo}_{+}(S)$ consisting of homeomorphisms whose support is outside of neighborhoods of stable maximal ends. Note that $\Gamma(S)$ is an open normal subgroup of $\operatorname{Homeo}_{+}(S)$.
@@
    A $2$-manifold $S$ is spacious if it is $S$ is weakly self-similar and for any simple closed curve $c$ in $S$, $\Gamma(S)\backslash\operatorname{Homeo}_{+}(S)\cdot c$ is finite.

### v3

@@ 1. Introduction
Let $S$ be an orientable surface. We say that $S$ is of finite type if its fundamental group is finitely generated. Otherwise, we will call $S$ is called a surface of infinite type. Equivalently, $S$ has infinite type if it has either infinitely many punctures, infinitely many genus, or both. A mapping class group of an infinite type surface $S$ is usually called a big mapping class group. Throughout this article we denote the big mapping class group of $S$ by $\mathrm{Map}(S)$. For a more detailed study of the big mapping class groups, we refer to the reader to [AV02].
@@ 1.2. First cohomology classes $H^{1}(\mathrm{FMap}(S),\mathbb{Z})$
The maximality of an end is derived from the partial order on the end space introduced in [MR23]. Mann and Rafi proved that under this partial order, the maximal element exists. Choose an immediate predecessor $z$ which both accumulates to maximal ends $A$ and $B$. (For precise definition, we refer Section 2.) We number each $z$ by integers in a sequence so that $\{z_{n}\}_{n\in\mathbb{Z}}$ accumulates to $A$ (or $B$) when $n$ goes to $\infty$ (or -$\infty$, respectively). The generalized shift $\eta_{A,B,z}$ is a shift map which sends $z$ of index $n$ to $z$ with index $n+1$. In simple terms, the cohomology class we investigate counts how many $z$’s are shifted. Collecting all such countings (including handle shifts), we have the following structure for $\mathrm{FMap}(S)$.
@@ 1.3. Normal generating set of big mapping class groups
There have been numerous studies on the (normal) generation of the big mapping class group, as finding a nice generating set for a given group is crucial for understanding its structure and properties. In [APV20], the authors prove that the set of Dehn twists topologically generates $\mathrm{Map}(S)$ if and only if $S$ has at most one non-planar end. Moreover, Patel and Vlamis [PV18] proved that if $S$ has at least $2$ ends accumulated by genus, then $\mathrm{PMap}(S)$ is topologically generated by Dehn twists and handle shifts. In [CC22], the authors show that for the finite type surface $S$ with a Cantor set $K$ embedded, $\mathrm{Map}(S-K)$ is normally generated by torsion normal generator of $\mathrm{Map}(S)$ unless $S$ has genus $2$ with $5k+4$ punctures for some $k\geq 0$. In [HHM+22], the authors studied topological properties of conjucagy class in $\mathrm{Map}(S)$. They showed that all conjugacy classes in $\mathrm{Map}(S)$ are meager and provided a condition for $\mathrm{Map}(S)$ to contain a dense conjugacy class. Additionally, as mentioned earlier, Vlamis ([VLA23], [VLA24]) proved that if $S$ is perfectly self-similar or telescoping, then $\mathrm{Map}(S)$ has a normal generator. Also, Mann and Rafi [MR23] give a necessary and sufficient condition for a tame surface $S$ to admit a coarsely bounded generating set of $\mathrm{Map}(S)$ and explicitly constructed the CB generating set.
@@
Also, let $C:=\max_{A,B\in\mathcal{M}(S)}|E_{cp}(A,B)|$. Then $\mathrm{Map}(S)$ is topologically normally generated by M(M+C1)M(M+C-1) number of generators.
@@ 1.4. Bounds of the rank of the abelianization of $\mathrm{Map}(S)$
    The commutator subgroup $\mathrm{Map}(S),\mathrm{Map}(S)]$ is Polish and both open and closed subgroup of $\mathrm{Map}(S)$.
@@
Corollary 8.10 Suppose $S$ is a finite connected sum of perfectly self-similar surfaces. Then the abelianization of $\mathrm{Map}(S)$ is finitely generated by at most M(M+C1)M(M+C-1).
@@ 2.2. Mann-Rafi’s classification
    $E(z)\cap(A-\{x_{A}\})\neq\emptyset\text{\qquad and \qquad}E(z)\cap(B-\{x_{B}\})\neq\emptyset$
@@ 3. Obstruction
$1\to\mathrm{PMap}(S)\xrightarrow{\iota}\mathrm{Map}(S)\xrightarrow{\pi}\operatorname{Homeo}(E(S),E^{G}(S))\to 1$
@@
[…]
@@ 4. Various shifts
$\mathcal{B}\cong\mathbb{R}\times[-1,1]-\left(\bigcup_{m\in\mathbb{Z}}D_{\frac{1}{2}}(m,0)\right)$
@@
We label $\partial_{m}$ to each boundary of the deleted disk centered at $(m,0)$. We also label $\partial_{\pm}$ as long boundaries, $\mathbb{R}\times\{\pm 1\}$.
@@
We index each $z$ so that $z_{k}$ accumulates to one maximal end $A$ as $k$ goes to -$\infty$, The other maximal end $B$ is accumulated by $z_{k}$’s with positive indices.
@@ 5. Cohomology classes of $\mathrm{FMap}(S)$
By construction, $\Phi_{z}(\eta_{A,B,z^{\prime}})=1$ if and only if $z\sim z^{\prime}$. In other words, $\Phi_{z}$ detects how many shifts of $z$ occur between $A,B$. This detection also works for other types of shift maps.(recall Definition 4.4 that there are $3$ types of generalized shifts.) Proposition 4.5 implies that even if the shift is not full, it differs only by the surface homeomorphism $T$ to the full shift. Since $T$ is generated by half twists (or $1/n$ Dehn twists), $\Phi_{z}(T)$ is equal to 0 by (3).
@@
Consider the standard basis $\{e_{i}\}$ of $\mathbb{Z}^{N-1}$, where $z\in E_{cp}(A_{0},A_{i})$ is an end corresponding to the $i$th entry. Let $\eta_{i}:=\eta_{A_{0},A_{i},z}$ . Then $\Theta(\eta_{i})=e_{i}$. Since all $\eta_{i}$’s commute each other, the sequence splits. ∎
@@ 6. From $\mathrm{FMap}(S)$ to $\mathrm{Map}(S)$
    $\mathcal{P}$ measures the parity of the permutation on $\{x_{1},\cdots,x_{n}\}$, which sends odd permutations to $1$ and even permutations to 0, induced by the forgetful map $\mathrm{Map}(S)\twoheadrightarrow S_{n}\twoheadrightarrow\mathbb{Z}/2\mathbb{Z}$.
@@
    Let $i,j$ be distinct integers in $\{1,\cdots,n\}$. Choose $z\in Im(x)$ and let $\eta$ be a shift map of $z$ from $x_{i}$ to $x_{j}$. If $Im(x)$ only contains a handle, then $x$ must be a type of ends which is accumulated only by genus and take the shift map $\eta$ as a handle shift from $x_{i}$ to $x_{j}$. Recall that $\Theta_{z}:\mathrm{FMap}(S)\to\mathbb{Z}^{n-1}$. Define $\widetilde{\Theta_{z}}:\mathrm{FMap}(S)\to\mathbb{Z}^{n-1}\to\mathbb{Z}\to\mathbb{Z}/2\mathbb{Z}$, where the second map is the sum of all entries in (n1)(n-1) tuple and the last map is a usual quotient map by $2\mathbb{Z}$.
@@
Let $f\in\mathrm{Map}(S)$. We first show that $\widetilde{\Theta_{z}}$ is invariant under conjugation. Conjugation by an elements in $\mathrm{FMap}(S)$ does not change the value, by its definition. Hence, the only conjugation which may invoke any differences is by half twists. Conjugation by half twists differs value to either permute entries in the (n1)(n-1) tuple, or change signs of some entries. Therefore, taking the sum and quotient by $2\mathbb{Z}$ is invariant.
@@ 7.1. Half-spaces
We first decompose the surface $S$ as follows. Choose a subsurface $K$ embedded in $S$ which is homeomorphic to a sphere with $4$ boundaries, such that SKS-K has $4$ connected components consisting of $S_{u}$ deleted with an open disk and $3$ homeomorphic copies of $S_{p}$ deleted with an open disk. By abusing the notation, we label each subsurface with $S_{u}$ and $S_{1},S_{2},S_{3}$, respectively. Note that the boundary of $S_{u}$ separates the $S_{u}$ part from the $S_{p}$ part. See Figure 7 also.
Figure 7. The surface decomposition by the 22-sphere with $4$ boundaries at the center. There is a $S_{u}$ on the top, and the other 33 subsurfaces are $S_{1},S_{2},S_{3}$, all homeomorphic to each other.
@@
Before we continue the argument, we first introduce which element has a dense conjugacy class. We follow the notation in [LV23b]. Let $S$ be uniquely self-similar surface and $x$ be a maximal end of $S$.
Proposition 7.8 (Proposition 4.2 in [LV23b]).
Let $c$ be a separating curve of $S$ and $\Omega_{c}$ denote the component of ScS-c such that $x\in E(\Omega_{s})$. Suppose $G$ be a subgroup of $\mathrm{Map}(S)$ as follows.
@@ 8. Normal generating sets
Let $G$ be a subgroup of $\mathrm{Map}(K,\partial K)$ whose elements induce the mapping class in $\mathrm{Map}(S)$, except the Dehn twists along the boundary of $K$. Then $G$ is normally generated by M(M2)M(M-2) number of elements.
@@
Suppose $\gamma$ is a separating curve that intersects $2$ boundaries of $L$ and one of the connected components of $K-\gamma$ is a pair of pants. See also Figure 8. The conjugacy class of $T_{\gamma}$ depends only on which boundaries $\gamma$ meets. This is because, if we have another $\gamma^{\prime}$, then there are appropriate half twists $g,g^{\prime}$ supported in $K-\text{int}(L)$ so that $gg^{\prime}(\gamma)=\gamma^{\prime}$. Therefore we only need $\frac{M(M-1)}{2}$ curves for each pair of connected components of $K-\text{int}(L)$ . Also if we choose another curve intersecting $2$ boundaries of $L$, then the Dehn twist along the curve can be generated by the lantern relation. Again by the lantern relation, the Dehn twist along any boundary of $L$ is generated by the half twists which are supported on the component of $K-\text{int}(L)$ which contains the boundary.
@@
Also, let $C:=\max_{A,B\in\mathcal{M}(S)}|E_{cp}(A,B)|$. Then $\mathrm{Map}(S)=\overline{\langle\langle\mathcal{S}\cup\mathcal{D}\cup\mathcal{H}\rangle\rangle}$. i.e., $\mathrm{Map}(S)$ is topologically normally generated by M(M+C1)M(M+C-1).
@@
Bounds on the minimal number of topological normal generators gives a bound of rank of its abelianization. By Field-Patel-Rasmussen [FPR22], they proved that the commutator subgroup of the mapping class group of surface SFS-F we questioned above is closed is finitely generated and discrete with the quotient topology.
@@
    The commutator subgroup $\mathrm{Map}(S),\mathrm{Map}(S)]$ is Polish and both open and closed subgroup of $\mathrm{Map}(S)$.
@@
Suppose $S$ is a finite connected sum of perfectly self-similar surfaces, possibly with finitely many punctures. Then the abelianization of $\mathrm{Map}(S)$ is finitely generated by at most M(M+C1)M(M+C-1).
@@ 9. Remarks on locally CB
In [LV23a], the authors proved that for an infinite type surface $S$, $\mathrm{Map}(S)$ has the virtual Rokhlin property (i.e., $\mathrm{Map}(S)$ contains a finite index subgroup which has the Rokhlin property) if and only if $S$ is spacious. More precisely, they found an open subgroup $\Gamma(S)$ which is an open, finite-index subgroup of $S$ with the Rokhlin property. We first recall the some definitions in [LV23a].
@@
    Suppose $S$ is a weakly self-similar and it is homeomorphic to $N$ or N#NN\#N for some uniquely self-similar surface $N$. Then $\mu\in E(S)$ is maximally stable if there is a nested subsequence $\Sigma_{n},n\in\mathbb{Z}_{\geq 0}$ such that $\Sigma_{n}$ is homeomorphic to $N$ with the open disk removed, $\Sigma_{n}$ is a neighborhood of $\mu$, and $\cap_{n\in\mathbb{Z}_{\geq 0}}$ is empty. Let $\Gamma(S)$ be a subgroup of $\operatorname{Homeo}_{+}(S)$ consisting of homeomorphisms whose support is outside of neighborhoods of stable maximal ends. Note that $\Gamma(S)$ is an open normal subgroup of $\operatorname{Homeo}_{+}(S)$.
@@
[…]